### v3

# Mask Image Watermarking

Abstract
We present MaskWM, a simple, efficient, and flexible framework for image watermarking. MaskWM has two variants: (1) MaskWM-D, which supports global watermark embedding, watermark localization, and local watermark extraction for applications such as tamper detection; (2) MaskWM-ED, which focuses on local watermark embedding and extraction, offering enhanced robustness in small regions to support fine-grined image protection. MaskWM-D builds on the classical encoder-distortion layer-decoder training paradigm. In MaskWM-D, we introduce a simple masking mechanism during the decoding stage that enables both global and local watermark extraction. During training, the decoder is guided by various types of masks applied to watermarked images before extraction, helping it learn to localize watermarks and extract them from the corresponding local areas. MaskWM-ED extends this design by incorporating the mask into the encoding stage as well, guiding the encoder to embed the watermark in designated local regions, which improves robustness under regional attacks. Extensive experiments show that MaskWM achieves state-of-the-art performance in global and local watermark extraction, watermark localization, and multi-watermark embedding. It outperforms all existing baselines, including the recent leading model WAM for local watermarking, while preserving high visual quality of the watermarked images. In addition, MaskWM is highly efficient and adaptable. It requires only 20 hours of training on a single A6000 GPU, achieving 15× computational efficiency compared to WAM. By simply adjusting the distortion layer, MaskWM can be quickly fine-tuned to meet varying robustness requirements.
1 Introduction
Image watermarking (potdar2005survey, ) is a crucial technique for embedding imperceptible information into images, serving purposes such as copyright protection, content authentication, and provenance tracking. With the proliferation of AI-generated content (AIGC) (rombach2022high, ; saharia2022photorealistic, ), the boundary between real and synthetic images has become increasingly blurred, making it especially important to develop robust watermarking schemes for content verification and traceability.
Traditional deep image watermarking methods (zhu2018hidden, ; tancik2020stegastamp, ; jia2021mbrs, ) typically perform global watermark embedding and extraction, treating the entire image as a uniform entity. However, these global approaches suffer from several critical limitations. 1. Lack of local watermark extraction: When an image undergoes heavy tampering, such as inpainting (yu2023inpaint, ; zhang2023adding, ), the watermark may survive only in a small, local region that remains untouched by manipulation. In such cases, global methods often fail to extract the watermark effectively. 2. Inability to localize the watermark: Even if a watermark is successfully extracted from the image, global methods cannot localize which region actually contains the watermark, making it difficult for fine-grained forensic analysis and fair judgment. 3. Lack of local watermark embedding: In scenarios where only specific regions of an image are valuable and need protection, or when different parts of the image originate from different sources and require distinct watermarking, global embedding is inherently incapable of providing the flexibility and granularity.
We argue that the training paradigm of traditional global watermarking methods, which treats the entire watermarked image as a whole for both encoding and decoding, prevents the encoder and decoder from developing spatial awareness. Specifically, the decoder cannot identify which regions of the image contain watermark signals for effective extraction, while the encoder lacks the ability to adaptively embed the watermark into specific spatial regions.
Based on the above analysis, we propose MaskWM, a simple, efficient, and flexible image watermarking framework. MaskWM introduces a masking mechanism during training to guide the model learn spatially aware embedding and extraction of local watermark signals. Depending on the stage which the mask is introduced, MaskWM offers two variants: MaskWM-D introduces the mask only during the decoding phase, enabling global watermark embedding while supporting local extraction. Specifically, by applying masks to retain only the selected regions of the watermarked images during extraction, the decoder is guided to identify which regions contain watermark signals and to focus on them for effective local extraction. MaskWM-ED introduces the mask during the encoding and decoding phases, allowing the embedding and extraction of the local watermark. In this setting, the encoder is trained to embed not only the watermark bits but also the spatial mask into the image. This allows the encoder to leverage the mask to adaptively allocate watermark strength to designated regions, while keeping the rest of the image nearly unaffected.
Extensive experiments demonstrate that MaskWM significantly outperforms existing baselines in both global and local watermark extraction, watermark localization, and local watermark embedding, while preserving image quality. Specifically, for local watermark extraction, MaskWM achieves a nearly 100% extraction accuracy even when only 5% of the image carries watermark signals. In terms of watermark localization, MaskWM demonstrates high precision in identifying watermark regions. Furthermore, although not explicitly trained for multi-watermark embedding, MaskWM maintains strong extraction and localization performance even when embedding up to 5 distinct watermarks in a single image. More importantly, MaskWM exhibits strong robustness across a wide range of distortions, including geometric distortions that typically break many existing watermarking methods. In addition to its effectiveness, MaskWM is highly efficient. Training the encoder-decoder model requires only approximately 20 hours on a single A6000 GPU, which is 15× less compute than the recent state-of-the-art local watermarking model WAM (sander2025watermark, ). MaskWM also scales effortlessly to different bit lengths (e.g., 32, 64, and 128), whereas WAM is inherently limited to 32-bit embedding and does not scale beyond that. Furthermore, MaskWM supports fast fine-tuning for different use cases. For example, MaskWM can reach a near 100% extraction accuracy against VAE-based adaptive attacks after just 20k training steps. These advantages make MaskWM a practical, efficient, and scalable solution for real-world applications.
2 Background
2.1 Image Watermarking
Image watermarking techniques can generally be categorized into two types: global watermarking and local watermarking methods. Global watermarking methods aim to extract watermark information from the entire image. Most traditional deep learning-based approaches fall into this category. These methods focus on achieving robustness against various types of distortions, ensuring that the embedded watermark can still be reliably recovered even when the image undergoes degradation. For example, MBRS (jia2021mbrs, ) specifically targets robustness against JPEG compression. Methods like StegaStamp (tancik2020stegastamp, ) and PIMoG (fang2022pimog, ) are designed to handle real-world physical distortions such as screen-shooting and print-shooting. More recent approaches like ZoDiac (zhang2024ZoDiac, ) and SuperMark (hu2024supermark, ) tackle adaptive attacks, while Robust-Wide (hu2024robust, ) and VINE (lu2024robust, ) focus on robustness against instruction-driven image editing.
In contrast, local watermarking methods focus on extracting watermark information from a specific region of the image. Recent methods, such as WAM (sander2025watermark, ) and our proposed MaskWM, belong to this category. WAM treats watermark extraction as a segmentation task (kirillov2023segment, ), predicting watermark bits at the pixel level and then averaging these per-pixel predictions to obtain the final result. While this fine-grained approach enables local watermark extraction, it also presents several challenges. 1. Limited extraction from small regions: When the watermarked area is very small, only a few pixels contribute to extraction, making naive averaging unreliable, especially under distortions. 2. Lack of scalability with longer messages: WAM struggles to scale beyond 32-bit messages, as training becomes increasingly difficult with longer bit lengths. 3. High computational cost: Training WAM is resource-intensive, requiring eight V100 GPUs for nearly a week, which limits its practicality. 4. Lack of native local embedding: WAM embeds watermarks globally and then crops for local focus, introducing embedding losses that reduce extraction robustness.
2.2 Watermark Localization
Watermark localization (zhang2024editguard, ; hu2025videoshield, ; sander2025watermark, ) refers to the ability to determine which regions of a watermarked image still contain watermark information after modifications. This capability enables the identification of unaltered content, serving as an active detection mechanism for tamper localization. Currently, image watermarking techniques primarily adopt two paradigms for watermark localization. The first paradigm embeds a one-dimensional copyright watermark alongside a two-dimensional localization watermark in the original image. During extraction, localization is based on the fragility of the localization watermark, which cannot be fully recovered from a modified image. Key methods in this category include EditGuard (zhang2024editguard, ) and OmniGuard (zhang2024omniguard, ). EditGuard embeds a solid-color template within the host image and attempts to recover it from a modified version. The difference between the recovered and original templates is calculated at each pixel, and a threshold-based decision identifies watermark-preserved regions. OmniGuard improves upon EditGuard by embedding a natural image as the template, enhancing fidelity. It also introduces a Degradation-aware Tamper Extractor, improving robustness in detecting tampered regions under distortion. This paradigm requires parallel extraction of both copyright and localization watermarks, which may affect image quality. Moreover, both watermarks need independent robustness, and the presence of the template watermark does not guarantee the presence of the copyright watermark.
The second paradigm, in contrast, embeds only a one-dimensional copyright bitstream and directly determines the presence or absence of watermark information at each pixel to achieve localization. Methods such as WAM (sander2025watermark, ) and our proposed MaskWM fall under this category. WAM employs a decoder that simultaneously performs pixel-wise watermark presence detection and copyright bit extraction. In contrast, our MaskWM incorporates a dedicated localization module within the decoder, focusing solely on watermark presence detection at each pixel. This approach is more lightweight and easier to optimize. Compared to the first watermark localization paradigm, this method ensures that the local watermarked regions strictly correspond to the areas containing copyright watermark information, enhancing interpretability. Additionally, it guarantees both the robustness of copyright watermark extraction and the robustness of localization.
3 Methodology
3.1 Design Principles
In general, we have three main objectives: local watermark extraction, watermark localization, and local watermark embedding. Among these, our primary goal is local watermark extraction, which aims to recover the embedded message from images where only a small, spatially local region contains the watermark signal. In practice, we find that achieving high performance on this task naturally necessitates solving the other two problems as well. We now identify three key reasons why traditional watermark models fail under this setting.
First, since the decoder is trained exclusively on globally watermarked images and has never encountered cases with only locally embedded watermarks, it fails to perform zero-shot extraction on such inputs. Second, the non-watermarked portions of an image interfere with the decoder’s extraction process, especially when the watermark occupies a small area, leading to extraction failure. Third, because the decoder is optimized for global watermark extraction, the encoder tends to dilute the watermark’s intensity over the entire image. This results in local regions having either insufficient or fragmented watermark strength, thereby exacerbating the extraction challenge.
To address these challenges, we propose MaskWM-D, which introduces a basic mask mechanism during the decoding stage to guide the decoder in identifying and focusing on watermark-containing regions. To solve the first issue, we retain the watermark only in the masked regions and set other regions’ pixels to zero, training the decoder to extract watermarks from partially watermarked images. For the second issue, we replace the non-masked regions with the original clean image and add a watermark localization module in the decoder to differentiate between watermarked and non-watermarked areas, reducing interference from irrelevant content. These two strategies directly address the first two issues by enabling the decoder to extract watermarks from local regions, while also indirectly mitigating the third issue by guiding the encoder to evenly distribute the watermark under the end-to-end training, thus facilitating the decoder’s extraction process.
To further enhance the encoder’s ability to address the third challenge, we propose MaskWM-ED, which incorporates the mask during the encoding stage to explicitly guide watermark placement. In MaskWM-ED, the mask is embedded into the image along with the watermark bits during training. This enables the encoder to learn to actively concentrate the watermark within the selected regions based on the embedded mask, thereby further improving the robustness of local watermarking.
3.2 Training
The overall end-to-end training pipeline is shown in Figure 1. It consists of four stages: (1) Mask Generation, (2) Watermark Embedding, (3) Watermark Masking, and (4) Watermark Extraction. In the following, we provide a detailed description of each stage.
Figure 1: The overall end-to-end training pipeline of MaskWM. (1) In the Mask Generation stage, we generate candidate masks from four predefined types and randomly select one mask $M$ for the subsequent stages. (2) In the Watermark Embedding stage, the encoder $\mathcal{E}$ embeds the watermark bits $W_{gt}$ into the original image $I_{orig}$, optionally using the mask $M$ (for MaskWM-ED), to produce the watermarked image $I_{wm}$. (3) In the Watermark Masking stage, the mask $M$ is used to fuse $I_{orig}$ and $I_{wm}$ (see Eq. 2), resulting in the fused image $I_{fuse}$, which is then subjected to a randomly selected distortion from a predefined distortion pool, yielding the distorted image $I^{{}^{\prime}}_{fuse}$. The masked region is then cropped using $M$ to obtain $I_{mask}$. (4) In the Watermark Extraction stage, the decoder $\mathcal{D}$ extracts the predicted mask $M_{pd}$ from $I^{{}^{\prime}}_{fuse}$ and the predicted watermark bits $W_{pd}$ from $I_{mask}$.
Mask Generation.
The mask generation process constructs a pool of candidate masks with diverse types, from which one mask $M$ is randomly selected for each image during training. We follow a similar mask generation strategy as LaMa (suvorov2022resolution, ) for image inpainting, utilizing four distinct types of masks: Full Mask, Rectangle Mask, Irregular Mask, and Segment Mask. These masks can enhance the model’s ability to handle watermark embedding and extraction under diverse conditions from different perspectives: Full Mask enables the model to embed and extract watermarks across the entire image, serving as a fundamental capability. Rectangle Mask focuses on regularly shaped local regions, encouraging the model to operate within confined areas of varying sizes. Irregular Mask introduces complex, arbitrarily shaped regions to improve robustness in non-uniform contexts. Lastly, Segment Mask targets semantically meaningful areas by leveraging object masks from the MS-COCO dataset (ms-coco, ), helping the model generalize to real-world scenarios.
Watermark Embedding.
We describe the process of embedding both the watermark bits $W_{gt}$ and the optional mask $M$ into the original image $I_{orig}$. We first randomly sample binary watermark bits $W_{gt}\in{0,1}$ of length $l$, which are transformed into a feature map $f\in\mathbb{R}^{C_{f}\times H\times W}$ using a lightweight CNN, where $C_{f}$ is the number of channels in $f$, and $H$ and $W$ denote the height and width of $I_{orig}$, respectively. This CNN consists of a linear layer that maps $W_{gt}$ to a tensor of shape $(1,l,l)$, followed by bilinear interpolation to $(1,H,W)$ and several Conv-Norm-ReLU (CNR) blocks that produce the final feature map $f$.
To embed the watermark, we concatenate the original image $I_{orig}$ with the watermark feature $f$ along the channel dimension. For the MaskWM-D, this results in a tensor of shape $(3+C_{f},H,W)$, promoting global watermark embedding. For the MaskWM-ED, the optional mask $M$ is further concatenated, yielding a tensor of shape $(3+C_{f}+1,H,W)$. The mask guides the model to focus on the selected regions during training, enabling local embedding within those specified areas.
The concatenated tensor is then passed through a U-Net (kirillov2023segment, ) to generate an intermediate encoded image $I_{enc}$. To obtain the final watermarked image $I_{wm}$, we apply a Just-Noticeable-Difference (JND) module (wu2017enhanced, ), which modulates the embedding signal based on human visual sensitivity to enhance the perceptual quality:
$I_{wm}=I_{orig}+\mu\times\mathrm{JND}(I_{orig})\times(I_{enc}-I_{orig}),$ (1)
where $\mu$ is the JND scaling factor to control the watermark strength. We explore several strategies to improve the visual quality of the watermarked image and find that JND modulation consistently delivers the best performance (see Appendix D.1 for details).
Watermark Masking.
We describe how the mask $M$ is used to process the watermarked image $I_{wm}$ for subsequent mask prediction and watermark extraction by the decoder. First, we generate a fused image $I_{fuse}$ by combining $I_{wm}$ and the original image $I_{orig}$, where the unmasked regions are replaced with the corresponding pixels from $I_{orig}$:
$I_{fuse}=I_{wm}\odot M+I_{orig}\odot(1-M).$ (2)
Next, a randomly selected distortion from a predefined distortion pool is applied to $I_{fuse}$, producing an augmented image $I^{\prime}_{fuse}$. This step follows a common practice in traditional watermarking methods to improve robustness against various transformations. Finally, we use the mask $M$ once more to isolate the watermarked regions of $I^{\prime}_{fuse}$, setting all other pixels to zero to obtain the input $I_{mask}$ for watermark extraction:
$I_{mask}=I^{\prime}_{fuse}\odot M.$ (3)
Watermark Extraction.
We describe how the decoder $\mathcal{D}$ extracts the predicted mask from $I^{\prime}_{fuse}$ and recovers the watermark bits from $I_{mask}$. To achieve these two objectives, $\mathcal{D}$ consists of two dedicated modules: a U2-Net (qin2020u2, ) for mask prediction, and a U-Net (ronneberger2015u, ) followed by a CNN for watermark extraction. Specifically, the U2-Net takes $I^{\prime}_{fuse}$ as input and predicts a mask $M_{pd}$. Meanwhile, the U-Net processes $I_{mask}$ to produce an intermediate feature $f^{\prime}$ with shape $(C_{f},H,W)$, which is then passed through a CNN to obtain the predicted watermark bits $W_{pd}$. Unlike the CNN used in the encoder $\mathcal{E}$, the CNN in the decoder $\mathcal{D}$ first applies several Conv–Norm–ReLU layers to further extract features from $f^{\prime}$. The resulting features are then interpolated to a fixed shape of $(1,l,l)$ and subsequently transformed by linear layers into a bit sequence of length $l$, yielding the final watermark prediction $W_{pd}$.
Training Objectives.
For all loss functions, we use Mean Squared Error (MSE). Specifically, the encoder loss is defined as:
$\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{MSE}}(I_{wm},I_{orig}).$ (4)
Note that we impose constraints only in the pixel space, as we find that this setup, combined with JND modulation, already achieves high visual quality. Introducing constraints in the feature space or using GAN-based losses would negatively impact the overall performance, as discussed in Appendix D.1. The decoder loss is formulated as:
$\mathcal{L}_{\text{dec}}=$ $\ \mathcal{L}_{\text{MSE}}(W_{pd},W_{gt})+\alpha\mathcal{L}_{\text{MSE}}(M_{pd},M),$ (5)
where $\alpha$ is a factor controlling the weight of the mask loss. The overall objective function is:
$\mathcal{L}_{\text{total}}=\beta_{\text{enc}}\cdot\mathcal{L}_{\text{enc}}+\beta_{\text{dec}}\cdot\mathcal{L}_{\text{dec}},$ (6)
where $\beta_{\text{enc}}$ and $\beta_{\text{dec}}$ are the weights for the encoder and decoder losses, respectively. Compared to conventional watermarking methods, our approach introduces only a mask loss at the decoder stage. As a result, it retains a simple yet effective objective, making it easy to extend.
3.3 Inference
Figure 2: Watermark localization and extraction process in our decoder during inference.
The inference process consists of three main stages: watermark embedding, localization, and extraction. Embedding is performed by the encoder. In MaskWM-D, the encoder embeds the watermark bits across the entire image. In MaskWM-ED, users provide both the watermark bits and a mask, allowing the encoder to embed the watermark into specific regions. Localization and extraction are performed by the decoder. As shown in Figure 2, unlike conventional global extraction methods, our approach first uses the decoder’s localization module to identify watermark-containing regions. The rest of the image is masked out by setting non-watermarked areas to zero, reducing noise during extraction. The watermark is then recovered from the retained regions. As discussed in Sec. 4.5, this strategy improves extraction robustness, especially when only small regions remain watermarked.
3.4 Usage Scenarios
This section outlines the usage scenarios of MaskWM-D and MaskWM-ED, which are tailored to different protection requirements. MaskWM-D is designed for full-image protection. It enables reliable watermark extraction even when parts of the image are tampered with, making it suitable for scenarios that demand global content integrity and tamper detection, such as copyright enforcement or image provenance verification. MaskWM-ED, on the other hand, focuses on protecting specific regions of interest, such as faces, logos, or sensitive content. It allows targeted verification or tracing if these regions are reused or misappropriated, without introducing watermarks across the entire image. This strategy not only supports privacy-aware or content-specific protection, but also tends to provide better robustness within the marked regions.
4 Experiments
4.1 Implementation Details
Training.
For all experiments, we train MaskWM on 83k images from the MS-COCO 2014 training set (ms-coco, ) and the training details are provided in Appendix C.1.1.
Evaluation.
To ensure fair comparison, we fix the image resolution to $512\times 512$. For baseline methods that do not support this resolution, we follow the resolution scaling strategy from TrustMark (bui2023trustmark, ) to interpolate the watermark strength (see Appendix C.2.1), which has been shown to preserve watermarking performance. To ensure comparable visual fidelity across variants, we set the JND scaling factor $\mu$ to 1.3 for MaskWM-D and 1.75 for MaskWM-ED. For robustness evaluation, we separately assess valuemetric and geometric distortions. For valuemetric robustness, we randomly sample from a set of ten common distortions, including JPEG Compression, Gaussian Filter, Gaussian Noise, Median Filter, Salt&Pepper Noise, Resize, Brightness, Contrast, Hue, and Saturation. For geometric robustness, we randomly sample from three typical transformations: Rotation, Perspective, and Horizontal Flip. These distortions collectively cover the vast majority of real-world transformations that watermarked images are likely to encounter in practical scenarios. Detailed parameter settings for each distortion are provided in Appendix C.2.2. We also evaluate robustness under each specific distortion individually and the detailed results are presented in Appendix D.2.
4.2 Global and Local Watermarking Comparison
Settings.
We compare MaskWM with seven recent open-source watermarking methods, including both global (e.g., StegaStamp (tancik2020stegastamp, ), SepMark (wu2023sepmark, ), TrustMark (bui2023trustmark, ), EditGuard (zhang2024editguard, ), Robust-Wide (hu2024robust, ), VINE (lu2024robust, )) and local (WAM (sander2025watermark, )) approaches. We use the clean (EditGuard-C) and degraded (EditGuard-D) variants of EditGuard, and the robust version (VINE-R) of VINE. For global watermarking, we evaluate on 1k images from the MS-COCO 2014 validation set using PSNR, SSIM, and Bit Accuracy, where PSNR and SSIM measure the visual quality of watermarked images, and Bit Accuracy evaluates watermark extraction performance. For local watermarking, we use all 41k validation images and evaluate Bit Accuracy under different watermarked area ratios. By default, MaskWM embeds 32 bits for fair comparison with WAM, though MaskWM supports flexible bit lengths (see Sec. 4.5). See Appendix C.2.3 for more implementation details and evaluation settings.
Global Watermarking Results.
The global watermarking results for all methods are summarized in Table 1. First, both MaskWM-D and MaskWM-ED achieve high visual fidelity, with PSNR scores above 39.5 and SSIM scores exceeding 0.98. These results outperform WAM and are only marginally lower than TrustMark and Robust-Wide. More importantly, under this high-fidelity setting, both variants of MaskWM still maintain near 100% bit accuracy, even under various valuemetric and geometric distortions. This demonstrates significantly better robustness compared to both global and local watermarking baselines. Notably, geometric distortions, which often break existing global watermarking methods, are effectively handled by our robust and reliable MaskWM framework.
Table 1: Comparison with baseline watermarking methods in terms of global watermarking. The best and the second best results are highlighted in bold and underlined, respectively.
Distortions
Method Bit Length PSNR $\uparrow$ SSIM $\uparrow$ No Distortion $\uparrow$ Valuemetric $\uparrow$ Geometric $\uparrow$
Global Watermarking Methods
StegaStamp (tancik2020stegastamp, ) 100 28.87 0.9019 0.9990 0.9976 0.6646
SepMark (wu2023sepmark, ) 30 35.73 0.9876 0.9957 0.9643 0.5086
TrustMark (bui2023trustmark, ) 100 41.19 0.9922 0.9996 0.9955 0.7868
EditGuard-C (zhang2024editguard, ) 64 37.27 0.9332 0.9991 0.5482 0.4925
EditGuard-D (zhang2024editguard, ) 64 32.30 0.8199 0.9999 0.5444 0.4975
Robust-Wide (hu2024robust, ) 64 41.58 0.9923 1.0000 0.9944 0.4951
VINE (lu2024robust, ) 100 36.04 0.9874 0.9997 0.9986 0.5012
Local Watermarking Methods
WAM (sander2025watermark, ) 32 39.32 0.9791 1.0000 0.9986 0.8979
MaskWM-D (Ours) 32 39.55 0.9814 1.0000 1.0000 0.9998
MaskWM-ED (Ours) 32 39.52 0.9828 1.0000 1.0000 1.0000
Local Watermarking Results.
The local watermarking results for all methods are presented in Figure 3. Global watermarking methods suffer a significant drop in extraction accuracy as the watermarked region shrinks, revealing their weakness in local watermarking tasks. In contrast, both WAM and MaskWM maintain high accuracy even with small watermarked areas, and MaskWM consistently outperforms WAM, especially under distortions, demonstrating greater robustness. Between our variants, MaskWM-ED performs better when the watermarked area is small, with the performance gap narrowing as the watermarked area increases.
Figure 3: Watermark extraction performance of different methods under different ratios of watermarked areas. The intervals of ratios are: 1-5%, 5-10%, …, 95-99%, 99-100%. We select the average value for each interval’s ratios to stand for the interval (e.g., 3% for 1-5%).
Visualized Results.
The visualized watermark patterns embedded by MaskWM-D and MaskWM-ED are shown in Figure 15 in Appendix D.6.
4.3 Watermark Localization Comparison
Settings.
We further compare MaskWM with EditGuard (zhang2024editguard, ) and WAM (sander2025watermark, ), two methods with watermark localization capabilities. Localization performance is evaluated using the local watermarking dataset described in Sec. 4.2, with Intersection-over-Union (IoU) metrics computed between the predicted and ground-truth regions for both the watermarked and unwatermarked areas.
Results.
The localization results of different methods are presented in Figure 4. First, MaskWM consistently achieves the best localization performance across nearly all watermark ratios and distortion conditions, with WAM showing a noticeable performance gap and EditGuard performing significantly worse. Second, MaskWM-ED outperforms MaskWM-D in localizing small regions, whether watermarked or unwatermarked, especially under distortion conditions. Interestingly, EditGuard-C shows an unusual rise in localization performance for unwatermarked regions as the watermark area increases under non-distorted conditions, a behavior not observed in EditGuard-D, possibly due to overfitting from being trained solely on clean data. The visualized localization results of different methods are provided in Appendix D.7.
Figure 4: Localization performance of different methods under different ratios of watermarked areas.
4.4 Performance Comparisons of Embedding Multiple Watermarks
Settings.
We further compare MaskWM-ED with WAM (sander2025watermark, ) under a multi-watermark setting by embedding up to five distinct watermarks into separate masked regions. The evaluation considers extraction accuracy and localization performance under a stricter masking constraint (5% area per region). We report the average watermark extraction accuracy across all embedded watermarks and the mean IoU of the predicted watermark regions. Detailed setup is provided in Appendix C.2.4.
Results.
Figure 5 presents the comparison results. First, our MaskWM-ED consistently outperforms WAM in both watermark extraction and localization, across all tested numbers of embedded watermarks and under various distortion conditions. Second, despite being trained solely with single-watermark supervision, MaskWM-ED generalizes well to multi-watermark settings, demonstrating strong scalability. However, under geometric distortions, MaskWM-ED’s extraction accuracy degrades as the number of embedded watermarks increases. This is mainly due to spatial transformations (particularly rotation) shrinking the watermark regions in the image corners, reducing the effective area available for extraction (see Figure 20 in Appendix D.8).
Figure 5: Performance comparison of watermark extraction and localization when embedding multiple watermarks. The bar colors transition from light to dark from left to right, representing the embedding of 1 to 5 different bit strings in a single image.
4.5 More Analysis
Scalability to Different Watermark Bit Lengths.
While WAM is constrained to 32-bit watermarks, our MaskWM readily scales to longer bit lengths. Detailed results are provided in Figure 12 in Appendix D.3. MaskWM maintains high extraction accuracy across various bit lengths. Even at 64 bits, the accuracy experiences only a slight drop and still outperforms WAM at 32 bits. While the accuracy decline becomes more noticeable at 128 bits, MaskWM consistently surpasses WAM under both no distortion and valuemetric distortion conditions. The only exception arises under geometric distortion, where a performance gap emerges when the watermarked region covers between 5% and 75%. These results underscore the scalability and optimization-friendly design of our method, emphasizing its practical advantages.
Enhancing Robustness against Adaptive Attacks via Fast Fine-tuning.
While MaskWM is trained on a broad range of common distortions, it is impractical to cover all possible scenarios. Fortunately, our framework supports task-specific fine-tuning, allowing it to adapt to new and emerging threats. As a demonstration, we consider adaptive attacks based on variational autoencoders (VAEs) (kingma2014auto, ), which reconstruct images and may inadvertently remove watermark signals. To counter this, we expand the distortion pool during fine-tuning by incorporating VAE modules from Stable Diffusion v1-4 (rombach2022high, ), Bmshj18 (ballé2018variational, ), and Cheng20 (cheng2020learned, ). Training details are provided in Appendix C.1.1. Fine-tuning for 20k steps on a single A6000 GPU ($\sim$5 hours) significantly improves robustness against these attacks, as shown in Figure 13 in Appendix D.4, outperforming existing baselines. While there is a slight drop in robustness to other distortions, MaskWM still maintains strong overall performance. These results demonstrate the adaptability of our approach to specific threats.
Importance of Localization before Extracting Local Watermarks.
We evaluate watermark extraction performance using three masking strategies when processing the watermarked image for extraction: (1) a full mask, which uses the entire image; (2) a predicted mask, which focuses on regions predicted to contain the watermark; and (3) a ground-truth mask, which focuses on the actual watermark-embedded regions. As shown in Figure 14 in Appendix D.5, when watermarks are embedded in small regions, using the full mask significantly reduces accuracy. In contrast, the predicted and ground-truth masks progressively improve performance. This highlights the importance of localizing watermarked regions before extraction to reduce interference from irrelevant content.
Computation Overhead Evaluation.
[…]
Table 2: Training and inference costs of different methods. The inference time is evaluated on a single A6000 with a batch size of 1 by averaging the total processing time over 1000 images.
# Params (M) Train Inference
Method Enc Dec Total Steps Batch GPU Time TFLOPs Memory Enc Time Dec Time Total Time Memory
EditGuard (zhang2024editguard, ) 3.6 2.6 6.2 250K 4 - - - - 0.074 s 0.080 s 0.154 s 2.15 GB
WAM (sander2025watermark, ) 1.1 96.0 97.1 3,680K 16 8 V100 90 h $1.13\times 10^{16}$ - 0.015 s 0.017 s 0.032 s 2.33 GB
MaskWM (Ours) 31.1 32.2 63.3 100K 16 1 A6000 20 h $7.74\times 10^{14}$ 25.84 GB 0.009 s 0.022 s 0.031 s 2.21 GB
5 Conclusion
In this paper, we propose MaskWM, a simple and efficient framework for both global and local image watermarking. Its core design introduces masks in encoding and decoding stages to guide the encoder and decoder to learn local embedding and extraction. Extensive experiments demonstrate MaskWM ’s superior performance in local watermark extraction and localization, along with high efficiency and adaptability. We hope our simple design can inspire future research to further enhance the practicality and functionality of watermarking models.
Acknowledgments
This research / project is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, and Infocomm Media Development Authority.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction are consistent with the actual contributions of the paper. They accurately summarize the proposed method, its motivations, and the key experimental findings presented in the main text.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of MaskWM in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides detailed descriptions of the experimental setup, including model architectures, training procedures, hyperparameters, evaluation metrics, and baselines. This information is sufficient to reproduce the main experimental results and verify the paper’s core claims, even without access to the code or data.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Both the code and data have been publicly released and detailed instructions are provided to enable faithful reproduction of the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The training and test details are provided in Sec. 4.1 and Appendix C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The main evaluation metrics (PSNR, SSIM, and Bit Accuracy) exhibit negligible variance across multiple runs in our setting, as they are deterministic or nearly deterministic given fixed seeds and models. As such, we did not include error bars, since their inclusion would not significantly affect the interpretation of the results or the validity of the main claims.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The type of compute workers, memory and time of execution are reported in Table 2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conforms with NeurIPS Code of Ethics and we provide impact statement in Appendix B.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We provide impact statement in Appendix B.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cited existing datasets and pre-trained models we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLMs were used as part of the core methodology or experimental pipeline. Any use of LLMs was limited to writing assistance and did not impact the scientific contributions or originality of the research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Limitations
Despite the relatively high PSNR and SSIM of the watermarked images produced by our method, noticeable artifacts are still present in some cases. To address this issue, we explored multiple approaches and ultimately adopted a JND-based method (wu2017enhanced, ) to effectively suppress artifacts (as shown in Appendix D.1). Nevertheless, some images still exhibit visible artifacts after watermark embedding, particularly in smooth background regions. This issue is also present in WAM (sander2025watermark, ) and reflects a common limitation in local watermarking methods, possibly arising from the inherent trade-off between imperceptibility and local robustness. Our focus in this work is to propose a watermarking framework centered on a mask mechanism, rather than to exhaustively pursue visual quality improvements. We have not further explored advanced strategies such as improved loss functions, auxiliary modules for visual enhancement, or more effective algorithms for watermark strength modulation. These aspects are left for future work to enhance the practicality and visual fidelity of our method.
Appendix B Impact Statement
We propose a simple, efficient, and flexible image watermarking framework whose core component, a masking mechanism, can be easily integrated into existing watermarking models. This mechanism enhances both the functionality and robustness of watermarking systems, significantly improving their practicality in real-world scenarios. The support for local watermark extraction allows reliable recovery of watermark information even when large portions of the image are tampered with. This is particularly valuable for forensic verification in areas such as misinformation mitigation and evidence integrity analysis. The ability to localize watermark regions enables systems to identify precisely which parts of an image contain embedded signals, facilitating transparent content attribution and enabling fair decision-making in legal and copyright-related contexts. Furthermore, local watermark embedding allows protection of specific regions within an image, which is useful when only certain parts require ownership marking or when content from multiple sources needs independent tracking and licensing. By improving spatial controllability while maintaining low computational cost and compatibility with existing architectures, our approach makes robust and fine-grained watermarking more accessible and deployable across diverse applications.
Appendix C More Details
C.1 Training
C.1.1 Details
All images are resized and center-cropped to $256\times 256$ during training. Training is conducted for 100k steps with a batch size of 16 on a single NVIDIA A6000 GPU. We use the AdamW optimizer with a learning rate of $1\times 10^{-4}$, and apply a cosine learning rate scheduler with 2k warm-up steps. We adopt an easy-to-hard training strategy inspired by TrustMark (bui2023trustmark, ). During the first 0.5k steps, the mask is set to full (i.e., all ones) and no distortion is applied. From step 0.5k to 1k, we introduce all types of masks. After 1k steps, distortions are added. The encoder loss weight $\beta_{\text{enc}}$ is fixed at 1, while the decoder loss weight $\beta_{\text{dec}}$ is initially set to 20 and linearly decayed to 0.2 over the first 5k steps. The mask loss weight $\alpha$ is set to 0.5. The JND module in the encoder is introduced and tuned starting from step 5k, with the scaling factor $\mu$ set to 1.
During fine-tuning with adaptive attacks, VAE-based distortions are applied with a 50% probability, while the original distortion types are retained for the remaining 50%. The hyperparameters are set as follows: $\beta_{\text{dec}}=0.3$, learning rate $1\times 10^{-4}$, and the quality levels for Bmshj18 and Cheng20 are both set to 5.
C.1.2 Distortions
Valuemetric Distortions.
During training, valuemetric robustness is enhanced by randomly sampling from a set of ten common distortions: JPEG Compression, Gaussian Filter, Gaussian Noise, Median Filter, Salt&Pepper Noise, Resize, Brightness, Contrast, Hue, and Saturation. The distortion parameters are set as follows:
  • JPEG Compression: quality factor = 50.
  • Gaussian Filter: kernel size = 1, sigma = 5.
  • Gaussian Noise: mean = 0, standard deviation = 0.1.
  • Median Filter: kernel size = 5.
  • Salt&Pepper Noise: noise ratio = 0.1.
Geometric Distortions.
To improve geometric robustness, we randomly sample from three typical geometric transformations: Rotation, Perspective, and Horizontal Flip. The specific configurations are:
  • Rotation: angle sampled from $90^{\circ},90^{\circ}]$.
  • Perspective: distortion scale sampled from $[0.1,0.5]$.
  • Horizontal Flip: no parameters.
C.2 Evaluation
C.2.1 Resolution Scaling
Algorithm 1 is adapted from the resolution scaling method described in TrustMark (bui2023trustmark, ). This algorithm enables a watermark model trained on images with a fixed resolution to be applied at arbitrary resolutions without sacrificing performance.
Input: Original image $\mathbf{x}$, [binary watermark vector $\mathbf{w}$]
Output: Watermarked image $\mathbf{y}$
Data: Embedding network $\mathbf{E}$ trained on the resolution of $m\times n$
1
[…]
$\mathbf{x}\leftarrow\mathbf{x}/127.5-1$ ;
// Normalize to range [-1, 1]
3 $\bar{\mathbf{x}}\leftarrow\text{interpolate}(\mathbf{x},\ (m,n))$
$\mathbf{r}\leftarrow\mathbf{E}(\bar{\mathbf{x}},\mathbf{w})-\bar{\mathbf{x}}$ ;
// residual image
4 $\mathbf{r}\leftarrow\text{interpolate}(\mathbf{r},\ (\text{H},\ \text{W}))$
5 $\mathbf{y}\leftarrow\text{clamp}(\mathbf{x}+\mathbf{r},\ -1,\ 1)$
$\mathbf{y}\leftarrow\mathbf{y}\times 127.5+127.5$
Algorithm 1 Resolution scaling - watermark embedding on arbitrary resolution images
C.2.2 Distortions
Valuemetric Distortions.
We apply ten types of valuemetric distortion with the following parameter settings to evaluate robustness:
  • JPEG Compression: quality factor = 60.
  • Gaussian Filter: kernel size = 1, sigma = 3.
  • Gaussian Noise: mean = 0, standard deviation = 0.05.
  • Median Filter: kernel size = 3.
  • Salt&Pepper Noise: noise ratio = 0.05.
  • Resize: scaling factor = 0.5.
  • Brightness Adjustment: range $(0.7,1.3)$.
  • Contrast Adjustment: range $(0.7,1.3)$.
  • Hue Adjustment: range (0.1,0.1)(-0.1,0.1).
  • Saturation Adjustment: range $(0.7,1.3)$.
Geometric Distortions
We apply three types of geometric distortion with the following parameter settings to evaluate robustness:
  • Rotation: angle sampled from $30^{\circ},30^{\circ}]$.
  • Perspective: distortion scale sampled from $[0.1,0.3]$.
  • Horizontal Flip: no parameters.
C.2.3 Global and Local Watermarking Comparison
Evaluation Setup.
For global watermarking, we sample 1,000 images from the MS-COCO 2014 validation set. We report PSNR and SSIM to assess visual quality, and Bit Accuracy to measure watermark extraction performance. For local watermarking, we construct a comprehensive evaluation set from all 41,000 images in the validation split. We divide the dataset into 12 subsets based on the ratio of masked area to the full image: 1–5%, 5–10%, 10–20%, …, 80–90%, 90–95%, and 95–99%. From each subset, 400 images are randomly selected. To simulate both inpainting and outpainting scenarios, we also include inverted masks, i.e., if the original mask covers aabb% of the image, the inverted mask covers (100b)(100{-}b)(100a)(100{-}a)%. This yields 800 image-mask pairs per subset and a total of 9,600 for evaluation.
Embedding Strategy.
Baselines and MaskWM-D embed watermarks across the entire image, but the unmasked regions are replaced with the original image to localize the watermark. MaskWM-ED embeds watermark bits only within the masked region and similarly restores the unmasked parts, ensuring true region-specific watermarking.
C.2.4 Multi-Watermark Embedding Setup
Figure 6: Checkerboard-like arrangement of masked regions for multi-watermark embedding. Different colors indicate regions where different watermarks bits are embedded, while black regions contain no embedded information.
Figure 6 illustrates the spatial arrangement of the five non-overlapping masked regions used in our multi-watermark experiments. These regions are fixed at the center, top-left, top-right, bottom-left, and bottom-right of the image, forming a checkerboard-like layout when all five are active. Unlike WAM, which allocates 10% of the image area to each region, we restrict each region to 5%, making the watermark extraction task more challenging. For each configuration (1 to 5 watermarks), we randomly sample 400 images from the MS-COCO 2014 validation set. During evaluation, we use OpenCV’s cv2.connectedComponents to segment the predicted mask into disjoint components, enabling region-wise watermark extraction and scoring.
Appendix D More Results
D.1 Effects of Different Visual Quality Enhancement Methods
Our experiments demonstrate that using only an MSE loss in the pixel space to constrain the difference between the watermarked and original images (referred to as the Base version) already yields high PSNR and SSIM scores. However, despite the favorable metrics, the resulting images often contain visible artifacts that may compromise perceived visual quality. To address this issue, we investigate several enhancement strategies: (1) incorporating a GAN (isola2017image, ) loss, (2) adding a perceptual constraint via LPIPS (zhang2018unreasonable, ) loss in the feature space, and (3) modulating the watermark signal using a Just Noticeable Difference (JND) (wu2017enhanced, ) module. We conduct experiments on the MaskWM-D variant, but the findings similarly hold for MaskWM-ED.
Figure 7: Visualization results of watermark using different visual quality enhancement methods. Zoom in to see more details. The numbers in parentheses indicate the corresponding loss weights. For example, GAN (0.001) means the GAN loss is assigned a weight of 0.001 in the total loss function, i.e., $\mathcal{L}_{\text{total}}=\beta_{\text{enc}}\mathcal{L}_{\text{enc}}+\beta_{\text{dec}}\mathcal{L}_{\text{dec}}+0.001\mathcal{L}_{\text{GAN}}$, and LPIPS (0.01) indicates that the LPIPS loss is weighted by 0.01, i.e., $\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{MSE}}(I_{wm},I_{orig})+0.01\mathcal{L}_{\text{LPIPS}}$. To ensure a fair comparison, we adjust either the strength of the added watermark residual or the JND modulation coefficient to maintain comparable PSNR and SSIM across different settings.
Figure 8: Local watermark extraction and localization performance of MaskWM-D trained with different visual quality enhancement methods.
As shown in Figure 7, when the weights of the GAN and LPIPS losses are low, they fail to effectively suppress artifacts. Increasing these weights reduces artifacts but adversely affects the performance of local watermark extraction and localization, as illustrated in Figure 8. In contrast, applying JND-based modulation proves more effective: it significantly reduces visible artifacts while maintaining performance comparable to the Base version. These findings suggest that, unlike global perceptual losses such as GAN or LPIPS, JND offers a more adaptive and content-aware modulation strategy. It effectively suppresses visual artifacts while preserving the watermark’s integrity, making it a practical choice for visual quality enhancement.
D.2 Results under Specific Distortions
See Figure 9, Figure 10, Figure 11.
Figure 9: Watermark extraction performance under various specific distortions.
Figure 10: Localization performance of the watermarked area under various specific distortions.
Figure 11: Localization performance of the unwatermarked area under various specific distortions.
D.3 Scalability to Different Watermark Bit Lengths
See Figure 12.
Figure 12: Watermark extraction performance of MaskWM-D and MaskWM-ED with different bits length. We also show the results of WAM for comparison.
D.4 Enhancing Robustness against Adaptive Attacks via Fast Fine-tuning
See Figure 13.
Figure 13: The effect of VAE fine-tuning on the robustness of MaskWM. Fine-tuning the VAE enhances robustness against VAE attacks, with minimal impact on the original robustness performance.
D.5 Importance of Localization before Extracting Local Watermarks
See Figure 14.
Figure 14: Watermark extraction performance using different masking strategies during decoding.
D.6 Visualization Results of Global and Local Watermark Embedding
See Figure 15.
Figure 15: Visualization results of global watermark embedding using MaskWM-D and local watermark embedding using MaskWM-ED. The residual image is acquired by $10\times|I_{wm}-I_{orig}|$ for observation, highlighting the residual more clearly. The same residual visualization strategy is applied in the following figures as well.
D.7 Visualization Results of Localization
See Figure 16, Figure 17, Figure 18, Figure 19.
Figure 16: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
Figure 17: Visualization results of watermark localization using different methods. The inpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
Figure 18: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
Figure 19: Visualization results of watermark localization using different methods. The outpainting results are obtained by applying stable-diffusion-2-inpainting (rombach2022high, ) to the masked regions for content reconstruction.
D.8 Visualization Results of Multiple Watermarks Embedding
See Figure 20.
Figure 20: Visualization of multi-watermark embedding and localization results. In the GT column, different colors in the mask indicate different watermark messages.